The difference between my son's age and my age is 28 years. Five years ago my age was double that of my son. Let my age now be X and my sons age now be Y.
Find the values of X and Y.

Answer by fractalier(6550)

Okay, let my age now be x and my sons age now be y. Then we write
x - y = 28 or x = y + 28
Five years ago
x-5 = 2(y-5)
Now substitute for x here and get
y+28 - 5 = 2(y-5)
y + 23 = 2y - 10
33 = y
and
then
x = 61
You are 61 and your son is 33.